# Two-dimensional nonlinear Schrödinger equations

### v1

@@ 3.1 Two-dimensional traveling wave type solutions with constant amplitude
Example 3. Let us consider a special case, by setting the linear relationship $g=af+b$ between the functions $f=f(r)$ and $g=g(r)$. In this case, the two-dimensional nonlinear Schrödinger equation (5) takes the form
@@
Solutions to this linear PDE can be found, for example, in [61]. Any particular solution $h=h(x,y)$ of the equation (29) generates an exact solution to the nonlinear PDE (28), which is described by the formulas (8) and (26) with $C_{1}=b$, where the function $r=r(x,y)$ is given implicitly by the relation $h(x,y)=rf(r)$.
In particular, equation (29) admits exact solutions
@@
Here $k$, $p$, and $q$ are free parameters. For $p>0$, $f>0$, and $f^{\prime}_{r}\geqslant 0$ it follows from (32) that $r\to 0$ as $z\to\pm\infty$, i.e. the resulting solution decreases exponentially rapidly for large $z$. This solution has a single maximum and describes a solitary wave, which is a soliton.

### v2

@@ 3.4 Time-periodic solutions with an amplitude depending on two spatial variables
Example 6. Let us consider in more detail a special case by establishing a linear dependence between the functions $f=f(r)$ and $g=g(r)$:
$g=af+b,$
where $a$ and $b$ are free parameters. In this case, the two-dimensional nonlinear Schrödinger equation (5) takes the form
@@
Solutions to this linear PDE can be found, for example, in [61]. Any particular solution $h=h(x,y)$ of the equation (43) generates an exact solution to the nonlinear PDE (42), which is described by the formulas (9) and (40) with $C_{1}=b$, where the function $r=r(x,y)$ is given implicitly by the relation $h(x,y)=rf(r)$. If conditions (16) hold, then this implicit relation defines a single-valued function $r=r(x,y)$.
In particular, equation (43) admits simple exact solutions
@@
Equation (43) also has the radially symmetric solution
$h=\begin{cases}AJ_{0}(\sqrt{a}\,\rho)+BY_{0}(\sqrt{a}\,\rho)&\text{if \ $a>0$},\\ AI_{0}(\sqrt{|a|}\,\rho)+BK_{0}(\sqrt{|a|}\,\rho)&\text{if \ $a<0$},\end{cases}\qquad\ \ \rho=\sqrt{x^{2}+y^{2}},$
where $A$ and $B$ are arbitrary constants, $J_{0}(z)$ and $Y_{0}(z)$ are the Bessel functions, and $I_{0}(z)$ and $K_{0}(z)$ are the modified Bessel functions.
More complex two-dimensional exact solutions of the linear PDE (43) in polar coordinates ($\rho$, $\theta$) have the form:
$h=\begin{cases}[A_{n}J_{n}(\mu\rho)+B_{n}Y_{n}(\mu\rho)][C_{n}\cos(n\theta)+D_{n}\sin(n\theta)]&\text{if \ $a=\mu^{2}>0$},\\ [A_{n}I_{n}(\mu\rho)+B_{n}K_{n}(\mu\rho)][C_{n}\cos(n\theta)+D_{n}\sin(n\theta)]&\text{if \ $a=-\mu^{2}<0$},\end{cases}$
where $n=1,\,2,\,\ldots\,$; $A_{n}$, $B_{n}$, $C_{n}$, and $D_{n}$ are arbitrary constants; $J_{n}(z)$ and $Y_{n}(z)$ are the Bessel functions; and $I_{n}(z)$ and $K_{n}(z)$ are the modified Bessel functions.
@@
Here $k$, $p$, and $q$ are free parameters. For $p>0$, $f>0$, and $f^{\prime}_{r}\geqslant 0$ it follows from (46) that $r\to 0$ as $z\to\pm\infty$, i.e. the resulting solution decreases exponentially rapidly for large $z$. This solution has a single maximum and describes a solitary wave.